# A Large Scale Analysis of Gender Biases in Text-to-Image Generative Models

### v1

@@ 6 Conclusion
We present the first large-scale analysis of gender bias in T2I models related to everyday activities and contexts. We generated 3,217,000 images (2,293,295 after filtering) for 3,217 prompts covering activities, contexts, objects, and occupations. Based on this extensive dataset, we conduct a detailed analysis of gender bias in the generated images. We find that models default to men in outputs, including for gender-neutral categories. Further, we find that models replicate and amplify gender norms and stereotypes; current models risk reinforcing and further solidifying harmful biases or limiting norms through widespread deployment. While models differ in the bias patterns they exhibit, likely also due to different training data, we do not observe that larger models are noticeably less biased than smaller or less capable models, suggesting that gender bias cannot be addressed by simply increasing the models’ capacity. Our research addresses a key gap in the literature on social bias in T2I models, significantly expanding and empirically grounding our understanding of gender stereotypes in synthetic images. Further research should incorporate intersectionality and explore how non-binary people are presented, including through prompts that explicitly evoke different gender identities and intersectional categories.
@@ Acknowledgements
Appendix A Prompts
In Sec. 3.1, we describe how we process prompts used to generate images and how we generate prompt variations. Here, we give further details on the precise prompts we use, and we also compare different LLMs to process prompts, justifying our choice of Yi-1.5-34B.
A.1 Prompt Processing LLM Prompts
@@
A.2 Prompt Processing LLM Comparison
Rewriting our prompts only requires shallow syntactical rewriting; therefore, we do not require particular reasoning skills from the LLM. Since we provide few-shot examples, we think most LLMs are suitable for our prompt processing. We decided to use Yi-1.5-34B due to its satisfactory performance. However, we compared four popular LLMs on 5 randomly sampled activities and found that all yielded the same results. In all cases, we used exactly the same prompts and system prompts. The compared LLMs are Yi-1.5-34B [98], Qwen2.5-32B-Instruct [97], Llama-3.1-8B-Instruct [27], and Phi-4 [1]. The results are in Tab. 3. All row-wise entries are identical, except for “drive around to look at sights with family in new home area” doesn’t insert the pronoun “their” before “family” in the processed prompt. We conclude that the choice of LLM is not crucial for our purposes, and we do not expect significant differences when using a different LLM than Yi-1.5-34B.
@@ A.2 Prompt Processing LLM Comparison
Table 3: Comparison of 4 different LLMs (Yi-1.5-34B [98], Qwen2.5-32B-Instruct [97], Llama-3.1-8B-Instruct [27], and Phi-4 [1]) on 5 randomly sampled activities. Prompts and system prompts are the same in all cases. All LLMs lead to the same processed prompts, suggesting the choice of LLM is irrelevant for our prompt processing purposes.
A.3 Prompt Clustering
HDBSCAN Settings. As described in Sec. 3.1, we use the HDBSCAN clustering algorithm to cluster prompt embeddings. Prompt embeddings are obtained from the all-mpnet-base-v2 model provided by [70] and reduced to 16 dimensions by UMAP [62]. For HDBSCAN, we use the implementation from scikit-learn with the following parameters:
@@ A.3 Prompt Clustering
Cluster summarization. We summarize prompt clusters using an LLM. An example of summarization is in Tab. 5. Concretely, we use Llama-3.3-70B-Instruct [27], with the following prompt:
@@
We decide to use Llama-3.3-70B-Instruct after comparing to other state-of-the-art LLMs, namely Qwen2.5-72B-Instruct [97], Yi-1.5-34B [98], and Phi-4 [1]. A comparison of the LLMs on 4 illustrative samples (activity clusters) is in Tab. 4. We notice that Llama-3.3-70B is superior in terms of how concise and fluent the resulting summaries are.
@@ 96 Baking activities Baking Sweet Activities Baking and Dessert Activities Baking and Baking Activities
shopping at walmart Grocery shopping
@@
Appendix B Image Generation
B.1 Diffusion Model Settings
@@
Generally, we use the hyperparameters (guidance scale, number of diffusion steps) recommended by the model authors. In all cases, the number of diffusion steps is 50, except for Flux-Schnell, where being a few-step-model [76] enables generating images with 4 diffusion steps. Guidance scales are as follows: 3.5 (Flux); 0.0 (Flux-Schnell); 3.5 (SD-3.5-Large); 4.5 (SD-3.5-Medium); and 7.0 (SD-3-Medium). Images are generated in $1024\times 1024$ for all models except Flux, where we generate images in $512\times 512$ for improved generation efficiency. After generation, all images are downscaled to $512\times 512$. Also note that, for Stable Diffusion models, we add the prompt prefix “a high-quality picture of” as we found this improves generation quality.
B.2 Prompt Following
@@
We use VQAScore [54] to measure how if generated images match their respective given prompts. VQAScore has been shown to yield better performance than related measures such as CLIPScore [41] or TIFA [47]. VQAScore uses an MLLM (clip-flant5-xxl which was trained by the authors of VQAScore specifically for this purpose) to predict the probability of answering “yes” when providing the MLLM the image and the following prompt:
@@
where “{prompt}” is replaced with the actual prompt used to generate the image. This yields a probability between 0 and 1, where a higher value indicates a stronger agreement between the prompt and the image. Therefore, higher values of VQAScore are desirable when generating images. However, models are expected to not generate good images for all prompts, and VQAScore is based on a statistical model with its own failure modes, introducing error-compounding effects. Also, for these purposes, it is not strictly necessary that models generate images that faithfully depict the prompt. We are interested in the associations of T2I models and gender, not general image quality. In Fig. 7, we show summary statistics of VQAScore values factored by combinations of models and prompt groups. We can see that in most cases, the VQAScore is above 0.5, indicating good prompt following.
Figure 7: Distributions of VQAScore values [54] factored by combinations of models and prompt groups. Higher scores are better.
B.3 Effect of Prompt Variations
@@
Appendix C Gender Indentification
C.1 MLLM Prompt
To identify perceived gender, we use the InternVL2-8B model. The InternVL2 model series [17, 18] was the strongest open model series when conducting experiments. We chose the 8B variant as it offers the best performance-efficiency tradeoff. Larger models do not perform better at perceived gender classification but incur a significant computational overhead.
@@ C.1 MLLM Prompt
Additionally, we randomly permute the option order (but not the letter order) to avoid label bias (e.g. the model preferring to predict the option letter “A”) [25].
We validate the performance of InternVL2-8B on VisoGender [38]. All images were labeled by human annotators for perceived gender. Specifically, we predict the gender of all 229 images in VisoGender that show a single person. 228 predictions are correct, meaning that perceived gender can be identified nearly perfectly by InternVL2-8B.
C.2 Nonbinary Gender Labels
We also evaluate if InternVL2-8B labels images as “nonbinary”. For this, we repeat the gender identification described in Sec. C.1, but add “nonbinary” as an option in addition to “female”, “male”, and “unclear/cannot tell”. Among the 5,675,715 person bounding boxes, only 332 receive the label “nonbinary”. This is not enough to conduct a meaningful quantitative analysis. However, we show 19 of 20 unique images generated by Flux that receive the “nonbinary” label. We removed one image that shows NSFW content. The images are in Fig. 9.
@@ C.2 Nonbinary Gender Labels
C.3 Images without Recognizable Gender
In Sec. 3.3, we filter images that show people but no person has a clearly recognizable gender according to InternVL2-8B. In total, 302,829 images are filtered by this criterion. One concern is that the gender of people in these images is perceived as nonbinary. Therefore, we inspect a sample of the filtered images but find that they are images where no gender cues are visible due to occlusion (shade, clothes), small size of people, or blurriness of people (in the background). Other images show only body parts, infants, or nonhuman creatures. We display 10 examples in Fig. 10.
@@ C.3 Images without Recognizable Gender
Appendix D Bias Agreement across Models
For each prompt group, we calculate the Spearman correlation between female ratios for all pairs of models. Correlations are only calculated on prompts that are not filtered for any model to ensure comparability. Results are in Fig. 11. We can see that all correlations are very high, especially for occupations and activities.
@@ Appendix D Bias Agreement across Models
Appendix E Detailed Analyses of Clusters
E.1 Retail contexts

### v2

@@ 5 Conclusion
We present a large-scale analysis of gender bias in T2I models, generating 3,217,000 images (2,293,295 after filtering) for 3,217 prompts covering activities, contexts, objects, and occupations. Across these, T2I models default to generating more images of men, including for gender-neutral prompts, confirming findings in prior work (Ghosh & Caliskan, 2023; Wu et al., 2024).
We consistently observe that scenarios with a high rate of female-gendered images portray women in traditional roles: as homemakers, while shopping, or engaged in arts and beauty in our activity prompts; as caring and service-oriented in our contexts and occupations; and with fashionable and soft objects. In contrast, men are associated with physical work, both in the household and at their jobs, working with machinery, and are strongly associated with business. While this reflects the greater numbers of women in caretaking roles and men in machinery-related or business roles that exist in society, our analysis shows that gender stereotypes are further amplified in T2I models.
While previous work could already prove bias amplification in occupations due to the existence of workforce labor statistics which do not exist for other scenarios (see Section I.5), we take a step further in analyzing bias amplification in activities by collecting statistics of a web-scale image-language corpus (LAION-400m), revealing that models can amplify bias beyond what is present in training data. These findings underscore the risk of reinforcing harmful norms through widespread deployment of T2I models.
To ensure validity, we filtered prompts and images for reliable gender evaluation. Although based on automatic methods, the strength of the patterns supports that they reflect spurious model biases. Our focus on binary gender is a limitation; we do not explore how identity-specific prompts (e.g., “female engineer”) might address or introduce stereotypes. Rather, our contribution is to analyze outputs for gender-neutral prompts to unpack underlying defaults and gender biases present in models. Future work should examine intersectionality and representation of non-binary identities. Additional limitations are discussed in Appendix J.
Ethics Statement
As text-to-image (T2I) models become widespread, assessing their societal impact, especially the reinforcement of social biases, is essential. We present a large-scale analysis of gender bias in five state-of-the-art T2I models, extending beyond occupation stereotypes to everyday situations, objects, and settings. The models frequently reproduce traditional gender roles, depicting women as homemakers and men in physical labor or business. These patterns raise ethical concerns. When used to synthesize training data for other systems, they can preserve existing inequalities. Repeated exposure to such images may also shape perceptions, perpetuating old stereotypes and creating new ones (Guilbeault et al., 2024).
By systematically measuring and characterizing these biases, we provide evidence to support more balanced image generation and to address biases in training data. We aim to inform researchers, developers, policymakers, and the public about subtle pathways through which AI can mirror inequality, and to help guide T2I development toward inclusive and fair practice.
However, our analysis is not without limitations: We automatically label perceived binary gender, excluding non-binary gender identities. This is mainly caused by technical limitations of current models, as discussed in Section 3.2. It is also important to note that assigning a person’s gender can be problematic because it cannot necessarily be perceived from an image, and gender is a spectrum. Therefore, good practice with images of real people is to have people self-identify their gender. However, T2I models create images that are not real, so assigning perceived gender to the images is more acceptable, as there is no risk of misidentifying a real person.
Reproducibility Statement
Images generated for this study, as well as the associated bounding boxes and perceived binary gender labels, are available on HuggingFace to enable further research on gender bias in T2I models: https://huggingface.co/datasets/LGirrbach/large-scale-gender-biases-analysis/. Additionally, we will make our code to reproduce experiments and analyses available on GitHub.
To improve clarity, the manuscript was polished for grammar and style using a large language model, with all final text reviewed and validated by the authors.
@@ Acknowledgements
Appendix A Broader Impact Statement
The growing use of T2I models makes it increasingly important to understand their potential effects on society, especially when it comes to reinforcing social biases. This study offers a large-scale analysis of gender bias in five leading T2I models, going beyond occupation-related stereotypes to uncover deeper patterns of bias in everyday situations, objects, and settings. Our results show that these models often reinforce traditional gender roles, such as frequently depicting women as homemakers and men in roles involving physical labor or business.
These biased images raise serious ethical concerns and can negatively affect many areas. When used to create training data for other machine learning applications, these biases may help preserve and spread existing inequalities. Likewise, repeated exposure to images showing traditional gender roles can shape how people view others, strengthening old stereotypes and possibly creating new ones (Guilbeault et al., 2024).
By carefully measuring and describing these gender biases, this research adds to the growing effort to create fairer AI systems. Our findings underline the urgent need to build more balanced image generation models and to address the biases in the data they are trained on. This work also aims to raise awareness among researchers, developers, policymakers, and the public about the quiet but widespread ways AI can mirror and amplify inequality. In the end, this study is a key step toward guiding AI development in a direction that is inclusive, fair, and better reflects the diversity of the world it serves.
Appendix B Compute Resources
Our experiments were conducted on an internal GPU cluster composed of a mix of NVIDIA A100 and NVIDIA H100 GPUs. Image generation required approximately 3,000 GPU-hours. Person bounding box detection and automatic perceived gender assignment took approximately 500 GPU-hours.
Appendix C Prompts
In Section 3.1, we describe how we process prompts used to generate images and how we generate prompt variations. Here, we give further details on the precise prompts we use, and we also compare different LLMs to process prompts, justifying our choice of Yi-1.5-34B.
C.1 Prompt Processing LLM Prompts
@@
C.2 Prompt Processing LLM Comparison
Rewriting our prompts only requires shallow syntactical rewriting; therefore, we do not require particular reasoning skills from the LLM. Since we provide few-shot examples, we think most LLMs are suitable for our prompt processing. We decided to use Yi-1.5-34B due to its satisfactory performance. However, we compared four popular LLMs on 5 randomly sampled activities and found that all yielded the same results. In all cases, we used exactly the same prompts and system prompts. The compared LLMs are Yi-1.5-34B (Young et al., 2024), Qwen2.5-32B-Instruct (Yang et al., 2024), Llama-3.1-8B-Instruct (Dubey et al., 2024), and Phi-4 (Abdin et al., 2024). The results are in Table 4. All row-wise entries are identical, except for “drive around to look at sights with family in new home area” doesn’t insert the pronoun “their” before “family” in the processed prompt. We conclude that the choice of LLM is not crucial for our purposes, and we do not expect significant differences when using a different LLM than Yi-1.5-34B.
@@ C.2 Prompt Processing LLM Comparison
Table 4: Comparison of 4 different LLMs (Yi-1.5-34B (Young et al., 2024), Qwen2.5-32B-Instruct (Yang et al., 2024), Llama-3.1-8B-Instruct (Dubey et al., 2024), and Phi-4 (Abdin et al., 2024)) on 5 randomly sampled activities. Prompts and system prompts are the same in all cases. All LLMs lead to the same processed prompts, suggesting the choice of LLM is irrelevant for our prompt processing purposes.
C.3 Prompt Clustering
HDBSCAN Settings. As described in Section 3.1, we use the HDBSCAN clustering algorithm to cluster prompt embeddings. Prompt embeddings are obtained from the all-mpnet-base-v2 model provided by Reimers & Gurevych (2019) and reduced to 16 dimensions by UMAP (McInnes et al., 2018). For HDBSCAN, we use the implementation from scikit-learn with the following parameters:
@@ C.3 Prompt Clustering
Cluster summarization. We summarize prompt clusters using an LLM. An example of summarization is in Table 6. Concretely, we use Llama-3.3-70B-Instruct (Dubey et al., 2024), with the following prompt:
@@
We decide to use Llama-3.3-70B-Instruct after comparing to other state-of-the-art LLMs, namely Qwen2.5-72B-Instruct Yang et al. (2024), Yi-1.5-34B Young et al. (2024), and Phi-4 Abdin et al. (2024). A comparison of the LLMs on 4 illustrative samples (activity clusters) is in Table 5. We notice that Llama-3.3-70B is superior in terms of how concise and fluent the resulting summaries are.
@@ 96 Baking activities Baking Sweet Activities Baking and Dessert Activities Baking and Baking Activities
shopping at walmart
Grocery
shopping
@@
Appendix D Image Generation
D.1 Necessity of Large Scale Image Generation
We generate 200 images per prompt (over 5 prompt variations), which is a very large number and requires significant computational scale. Therefore, we provide evidence of why it is necessary to generate so many images to gain the precise insights that we provide. Specifically, our large number of images per prompt is required for statistical precision. We conduct a subsampling experiment to quantify the stability of $R_{f}$ with a varying number of images. For each prompt, we create $1,000$ bootstrap samples for sample sizes ($n\in\{20,25,\ldots,200\}$) and calculate the mean absolute deviation of the sample $R_{f}$ from the $R_{f}$ calculated on our full data. The resulting mean absolute deviation values in the following table demonstrate that smaller sample sizes lead to considerable measurement error:
Model 20 25 30 40 50 100 150 200
Flux $0.060$ $0.052$ $0.050$ $0.044$ $0.039$ $0.029$ $0.023$ $0.022$
Flux-Schnell $0.061$ $0.055$ $0.051$ $0.045$ $0.041$ $0.031$ $0.026$ $0.025$
SD-3.5-Large $0.061$ $0.054$ $0.050$ $0.043$ $0.039$ $0.028$ $0.022$ $0.021$
SD-3.5-Medium $0.058$ $0.052$ $0.048$ $0.042$ $0.038$ $0.028$ $0.022$ $0.021$
SD-3-Medium $0.047$ $0.043$ $0.039$ $0.035$ $0.032$ $0.024$ $0.020$ $0.019$
Only with at least 100 images per prompt does the average deviation reliably drop below 3 percentage points. For smaller samples, the noise could easily obscure the effects we aim to measure. For instance, a 6 percentage point average deviation for a sample size of 20 is too high for precise analysis across thousands of prompts.
These findings strongly support the scale of our study. While smaller samples might be sufficient to merely show that some bias exists, they are inadequate for the goal of our work, which is to precisely quantify and compare gender biases across different models and a broad range of concepts.
D.2 Diffusion Model Settings
@@
Settings. Generally, we use the hyperparameters (guidance scale, number of diffusion steps) recommended by the model authors. In all cases, the number of diffusion steps is 50, except for Flux-Schnell, where being a few-step-model (Sauer et al., 2024) enables generating images with 4 diffusion steps. Guidance scales are as follows: 3.5 (Flux); 0.0 (Flux-Schnell); 3.5 (SD-3.5-Large); 4.5 (SD-3.5-Medium); and 7.0 (SD-3-Medium). Images are generated in $1024\times 1024$ for all models except Flux, where we generate images in $512\times 512$ for improved generation efficiency. After generation, all images are downscaled to $512\times 512$. Also note that, for Stable Diffusion models, we add the prompt prefix “a high-quality picture of” as we found this improves generation quality.
Ablation on Classifier-Free Guidance (CFG). The role of CFG in the quality-diversity trade-off makes it a potential candidate for influencing bias. To investigate this and make sure our work is not affected by it, we conduct an ablation study. We sample 50 “activity” prompts, stratified by their average female ratio, and re-generated images using two settings: (1) CFG turned off, and (2) a low CFG scale of 2.0.
Turning off CFG or lowering the guidance scale significantly degrades image quality for SD-3.5-Medium and SD-3-Medium, rendering most images unusable for analysis. We therefore proceed with the three models that produced coherent images (Flux, Flux-Schnell, SD-3.5-Large). The table below shows the mean absolute deviation of the female ratio $R_{f}$ from the values in our original study, where we use recommended CFG values for each model.
Setting Flux Flux-Schnell SD-3.5-Large
CFG off $0.077$ 0.003-0.003 $0.004$
Guidance Scale=2 $0.024$ 0.001-0.001 $0.004$
From this, we derive the following key findings: First, for models where CFG is a factor (Flux, SD-3.5-Large), changing the guidance scale has small or inconsistent effects on gender bias. Second, lowering or disabling CFG severely impacts the basic image quality of several models (SD-3.5-Medium, SD-3-Medium).
Therefore, we conclude that deviating from the recommended CFG settings introduces a strong confounding variable. Any observed changes in bias could be artifacts of the model’s failure to generate coherent subjects. Therefore, to ensure a fair and meaningful evaluation of bias on high-quality, in-distribution generations, we think that using the recommended guidance scale values is the most methodologically sound approach.
D.3 Prompt Following
@@
We use VQAScore (Lin et al., 2024) to measure how if generated images match their respective given prompts. VQAScore has been shown to yield better performance than related measures such as CLIPScore (Hessel et al., 2021) or TIFA (Hu et al., 2023). VQAScore uses an MLLM (clip-flant5-xxl which was trained by the authors of VQAScore specifically for this purpose) to predict the probability of answering “yes” when providing the MLLM the image and the following prompt:
@@
where “{prompt}” is replaced with the actual prompt used to generate the image. This yields a probability between 0 and 1, where a higher value indicates a stronger agreement between the prompt and the image. Therefore, higher values of VQAScore are desirable when generating images. However, models are expected not to generate good images for all prompts, and VQAScore is based on a statistical model with its own failure modes, introducing error-compounding effects. Also, for these purposes, it is not strictly necessary that models generate images that faithfully depict the prompt. We are interested in the associations of T2I models and gender, not general image quality. In Fig. 6, we show summary statistics of VQAScore values factored by combinations of models and prompt groups. We can see that in most cases, the VQAScore is above 0.5, indicating good prompt following.
Figure 6: Distributions of VQAScore values Lin et al. (2024) factored by combinations of models and prompt groups. Higher scores are better.
D.4 Effect of Prompt Variations
@@
Appendix E Gender Indentification
E.1 MLLM Prompt
To identify perceived gender, we use the InternVL2-8B model. The InternVL2 model series (Chen et al., 2024b; c) was the strongest open model series when conducting experiments. We chose the 8B variant as it offers the best performance-efficiency tradeoff. Larger models do not perform better at perceived gender classification but incur a significant computational overhead.
@@ E.1 MLLM Prompt
Additionally, we randomly permute the option order (but not the letter order) to avoid label bias (e.g. the model preferring to predict the option letter “A”) (Dominguez-Olmedo et al., 2024).
We validate the performance of InternVL2-8B on VisoGender (Hall et al., 2024). All images were labeled by human annotators for perceived gender. Specifically, we predict the gender of all 229 images in VisoGender that show a single person. 228 predictions are correct, meaning that perceived gender can be identified nearly perfectly by InternVL2-8B. Additionally, we evaluate InternVL2-8B on the SocialCounterfactuals dataset (Howard et al., 2024), which contains over 170,000 synthetic images from an older Stable Diffusion model. On this dataset, InternVL achieves 99.7% accuracy and a Cohen’s kappa of 0.994 against the ground truth labels, which means near-perfect performance on synthetic images with known attributes.
Finally, we also conducted a study on our own generated images. We randomly sampled 120 bounding boxes (40 labeled male, 40 female, 40 unclear by InternVL2-8B) and had them labeled by two human annotators unaffiliated with this submission. We then compared the MLLM labels to the human labels.
The model-to-human agreement (Annotator 1: $\kappa=0.78$) and Annotator 2: $\kappa=0.71$) is substantial and nearly identical to the human-to-human inter-annotator agreement ($\kappa=0.775$). Here, it is critical to note that InternVL2-8B performs on par with a human annotator for this task, as its agreement with a human is bounded by the agreement between two humans. This validates our use of InternVL2-8B for labeling perceived gender in our generated images.
E.2 Nonbinary Gender Labels
We also evaluate if InternVL2-8B labels images as “nonbinary”. For this, we repeat the gender identification described in Section E.1, but add “nonbinary” as an option in addition to “female”, “male”, and “unclear/cannot tell”. Among the 5,675,715 person bounding boxes, only 332 receive the label “nonbinary”. This is not enough to conduct a meaningful quantitative analysis. However, we show 19 of 20 unique images generated by Flux that receive the “nonbinary” label. We removed one image that shows NSFW content. The images are in Fig. 8.
@@ E.2 Nonbinary Gender Labels
E.3 Images without Recognizable Gender
In Section 3.3, we filter images that show people but no person has a clearly recognizable gender according to InternVL2-8B. In total, 302,829 images are filtered by this criterion. One concern is that the gender of people in these images is perceived as nonbinary. Therefore, we inspect a sample of the filtered images but find that they are images where no gender cues are visible due to occlusion (shade, clothes), small size of people, or blurriness of people (in the background). Other images show only body parts, infants, or nonhuman creatures. We display 10 examples in Fig. 9.
@@ E.3 Images without Recognizable Gender
Appendix F Comparison of Bias Across Models
F.1 Bias Agreement Across Models
For each prompt group, we calculate the Spearman correlation between female ratios for all pairs of models. Correlations are only calculated on prompts that are not filtered for any model to ensure comparability. Results are in Fig. 10. We can see that all correlations are very high, especially for occupations and activities. This means that models generally exhibit similar biases, although minor differences exist. The similarity of biases across models strengthens our conclusion that biases represented by the models included in this study will likely hold for other models as well.
@@ F.1 Bias Agreement Across Models
F.2 Comparison of Bias Strength across Models
Bias Direction. To see if any model consistently leans more male or female, we calculated the deviation of each model’s female ratio $R_{f}$ from the average $R_{f}$ across all models for each prompt. As shown in the table below, the results are mixed.
Model Activities Contexts Objects Occupations
Flux $0.031$ 0.045-0.045 $0.002$ 0.040-0.040
Flux-Schnell $0.032$ 0.078-0.078 0.014-0.014 0.028-0.028
SD-3.5-Large 0.004-0.004 $0.038$ $0.066$ 0.005-0.005
SD-3.5-Medium $0.001$ $0.089$ 0.028-0.028 $0.029$
SD-3-Medium 0.063-0.063 0.008-0.008 0.033-0.033 $0.045$
In a few cases, models significantly deviate from the model mean, for example SD-3.5-Large generates more women than the average for object prompts, and SD-3.5-Medium for context prompts. Conversely, Flux-Schnell generates fewer women for context prompts, and SD-3-Medium for activity prompts. However, no model consistently shows a positive or negative deviation across all categories.
Bias Intensity (Skew). Next, we measure the intensity of bias by calculating the entropy of the gender distribution for each prompt group. Lower entropy indicates a more skewed distribution (i.e., generations are heavily skewed towards one gender), signifying more intense bias.
Model Activities Contexts Objects Occupations
Flux $0.673$ $0.664$ $0.730$ $0.433$
Flux-Schnell $0.676$ $0.636$ $0.806$ $0.422$
SD-3-Medium $0.485$ $0.532$ $0.673$ $0.342$
SD-3.5-Large $0.618$ $0.730$ $0.802$ $0.440$
SD-3.5-Medium $0.591$ $0.733$ $0.756$ $0.418$
This analysis shows a clear and consistent trend. SD-3-Medium consistently produces the lowest entropy (most skewed) outputs, while SD-3.5-Large is the most balanced (highest entropy). The Flux models fall in between. This allows us to rank the models by the intensity of their gender bias: SD-3.5-Large (most balanced) >> Flux/Flux-Schnell >> SD-3.5-Medium >> SD-3-Medium (most skewed).
This ranking suggests that larger, more recent models may produce more balanced gender representations. However, this conclusion should be taken with caution, since information on training data and objectives is not public. Future work should investigate this further.
Bias vs. Image Diversity. We investigate the relation between gender bias in generated images and their diversity. For this, we use DINOv2 image embeddings and calculate the average pairwise cosine similarity for all images generated for a given prompt group. We use this as an inverse proxy for diversity, where lower scores indicate higher diversity. The results are summarized below:
Model Activities Contexts Objects Occupations
Flux $0.71$ $0.66$ $0.58$ $0.72$
Flux-Schnell $0.71$ $0.69$ $0.60$ $0.75$
SD-3.5-Large $0.72$ $0.71$ $0.61$ $0.75$
SD-3.5-Medium $0.73$ $0.72$ $0.61$ $0.77$
SD-3-Medium $0.76$ $0.76$ $0.63$ $0.79$
Combining these findings with our results on bias strength suggests that higher image diversity may not directly translate to less bias. Although models can be ranked clearly by their diversity and skew, as shown above, which are naturally related, this ranking does not correspond to a clear ranking by bias strength. This suggests that improving general capabilities like image diversity may not be a solution for mitigating bias.
Appendix G Why LAION-400M is a Suitable Baseline for Bias Amplification
In Section 4, we use context-gender cooccurrences in LAION-400M as a baseline to detect bias amplification. This raises the question of why LAION-400M is a suitable baseline. While we do not have access to the exact training data of the models, which is certainly larger and likely more diverse than LAION-400M, our use of LAION-400M as a proxy is justified and, we argue, makes our findings on bias amplification more robust. Our reasoning is twofold:
First, as noted by Udandarao et al. Udandarao et al. (2024), concept frequencies and scaling trends are consistent across different web-scale datasets. This suggests LAION-400M is a reasonable, publicly available proxy for the type of data these models are trained on. Second, LAION-400M is likely a conservative baseline. Assuming the premise is true, that the proprietary training sets are larger and more balanced than LAION-400M, it would mean our baseline is more biased than the models’ actual training data. In this scenario, observing that models still amplify bias relative to our LAION-400M baseline makes our conclusion even stronger. The models are not only failing to mitigate the bias in comparison to our proxy dataset, but they are also amplifying it.
Therefore, we argue our conclusions are reliable because LAION-400M serves as a reasonable and, importantly, conservative baseline for measuring bias amplification.
Appendix H Detailed Analyses of Clusters
H.1 Retail contexts